# Can LLMs handle WebShell detection? Overcoming Detection Challenges with Behavioral Function-Aware Framework

### v1

@@ 5 Related Work
Conclusion
@@ Appendix D Results
Table 4: Performance Comparison of BFAD-Enhanced Models Against Baselines. This table compares traditional ML and DL models with large and small-scale LLMs, both standalone and enhanced with BFAD.
Category Model Accuracy Precision Recall F1 Score
Sequence Baselines GloVe+SVM 96.20% 93.30% 94.30% 93.80%
CodeBERT+RF 96.30% 94.00% 95.60% 94.80%
Graph Baselines GCN 96.90% 94.40% 95.30% 94.90%
GAT 98.37% 99.52% 97.39% 98.87%
LLM Baselines (Large) GPT-4 97.27% 100.00% 85.98% 92.46%
LLaMA-3.1-70B 98.01% 97.31% 92.36% 94.77%
Qwen-2.5-Coder-14B 98.64% 99.32% 93.63% 96.39%
Qwen-2.5-3B 93.72% 78.03% 91.84% 84.37%
Qwen-2.5-1.5B 43.62% 34.61% 95.77% 50.84%
Qwen-2.5-0.5B 19.47% 18.65% 100.00% 31.44%
LLM + BFAD GPT-4 99.75% 100.00% 98.71% 99.35% (+6.89)
LLaMA-3.1-70B 99.38% 98.72% 98.09% 98.40% (+3.63)
Qwen-2.5-Coder-14B 98.76% 98.68% 94.90% 96.75% (+0.36)
Qwen-2.5-Coder-3B 78.89% 46.67% 100.00% 63.64% (+7.71)
Qwen-2.5-3B 97.39% 88.64% 99.36% 93.69% (+9.32)
Qwen-2.5-1.5B 80.40% 48.51% 100.00% 65.33% (+14.49)
Qwen-2.5-0.5B 91.94% 71.10% 98.73% 82.67% (+51.23)

### v2

@@ 5 Related Work
6 Limitations and Future Work
Our research highlights several limitations that present important directions for future investigation in this emerging field.
6.1 Dataset Generalization
A fundamental challenge in WebShell detection is the scarcity of real-world data due to privacy concerns. Current research, including this study, mainly relies on public datasets, which may not fully represent the diversity and sophistication of novel threats. This reliance makes it difficult to rigorously evaluate a model’s ability to generalize to unseen attack vectors. This limits rigorous evaluation of a model’s ability to generalize against novel and heavily obfuscated threats. To address this, future work should focus on using LLMs to generate synthetic benchmarks. A large-scale, LLM-generated dataset would provide a privacy-preserving and diverse alternative for stress-testing model robustness against a wider range of attack vectors.
6.2 Model Robustness and System Adaptability
While our BFAD framework is effective, the underlying static analysis approach is vulnerable to sophisticated obfuscation and zero-day threats. Enhancing the core model and building adaptive deployment systems are critical next steps. To improve model robustness, future work could integrate dynamic analysis, examining runtime function call traces from sandboxed execution to unmask behavior hidden by obfuscation. Additionally, since our analysis revealed semantic understanding bottlenecks even in large models, domain-specific fine-tuning of code-centric LLMs on curated WebShell datasets is a crucial step to deepen their expertise and improve detection accuracy. For practical and resilient deployment, these enhanced models could be integrated into advanced architectures. A multi-stage ”fast and slow” system could use a smaller LLM for rapid, efficient filtering, while reserving a larger reasoning model for in-depth analysis of high-risk samples. To ensure long-term effectiveness against evolving threats, this pipeline could be governed by an autonomous agent that continuously identifies new attack patterns and automatically updates the detection system.
7 Conclusion